# mini-vec2vec: Scaling Universal Geometry Alignment with Linear Transformations

### v3

@@ 4.5 Iterative Refinement
This refinement strategy proceeds by transforming embeddings from space $A$ to $B$ using $\mathbf{W}$, averaging their nearest neighbors in space $B$ based on cosine similarity, and obtaining a new orthogonal transformation with Procrustes analysis – similar to the process described in the Mapping Estimation stage (Section 4.4), with the crucial difference that now the similarity is computed in the ambient space $B$ rather than relative space. The method is cheap and it runs for many iterations (50–100), using subsampling to boost computational efficiency further.
@@
This strategy aims to improve the large-scale matching of embeddings by matching cluster centroids. We begin by clustering the embeddings in space $A$. We then transform the $A$ cluster centroids, and cluster the $B$ embeddings with the transformed $A$ centroids as initialization. The use of cluster centroids creates more stable pseudo-parallel pairs based on cluster assignments rather than individual nearest neighbors. The use of the transformed centroids as initialization follows from the assumption that they cannot be far off from cluster centroids in space $B$, so they will tend to converge to their counterpart centroids in space $B$, rather than arbitrary cluster centroids. This way, the seeded clustering will only correct minor errors in the transformation, and move the centroids to their “right positions” in space $B$. Similar to the first strategy, we use exponential smoothing.
Crucially, we find that one iteration of Refine-2 improves the alignment obtained by Refine-1. Two or more lead to a slight deterioration of the alignment, although this might be partly explained by our use of a number of clusters smaller than the dimensionality.
@@ 5 Experiments
We replicate the main experiment from jha2025vec2vec. We run our algorithm for all pairs of encoders used in jha2025vec2vec. Details about the encoders are provided in Table 2. Experiments use k-means clustering from scikit-learn (scikit-learn) and QAP is solved with the scipy (2020SciPy-NMeth) implementation of 2-OPT. We re-run the 2-OPT alignment algorithm 30 times and choose the permutation with the highest alignment score. This helps improve the robustness of the alignment, as a single 2-OPT is often not enough.
We take a random subset of the Natural Questions dataset (natural_questions) of size $60,000$ and compute sentence embeddings under all possible encoders. We leave $8192$ sentences for evaluation. For each pair, we split the remaining embeddings equally between the source encoder (space $A$) and target encoder (space $B$), so there is no overlap between the source and target sentences. All experiments are run on a Colab notebook with a CPU runtime and repeated three times for each pair to demonstrate consistency. We compare our results to the numbers reported in jha2025vec2vec.
@@
We have found that the method is extremely robust to hyperparameter choices, as long as they are in a reasonable range. We make an economical choice of hyperparameter values due to the large number of runs, but we find the choice to be sufficiently robust still. We use $s=30$ runs of relative representation alignment with $c=20$ clusters. In Mapping Estimation, we use $k=50$ neighbors. We set the sampling rate in Refine-1 to $n_{s}=10,000$ and utilize $k^{\prime}=50$ neighbors; In Refine-2, we use $c^{\prime}=500$ clusters. For exponential smoothing, we use $\alpha=0.5$. The full hyperparameter configuration used in the experiments is also documented in the code linked. For the $\textit{gtr}\leftrightarrow\textit{gte}$ pair, we increase the number of clusters to $c=30$. We find that for this pair, the algorithm often converges to a sub-optimal solution in the $c=20$ regime.
@@ Results
In Table 1, we demonstrate that our approach matches or exceeds vec2vec across (almost) the entire table. The improvement is most noticeable in the rank metric. Importantly, the method is extremely robust as is evidenced by the small standard deviation, already by the end of the first refinement step. Unlike vec2vec, our method does not collapse on the $\textit{e5}\leftrightarrow\textit{gtr}$ pair or the $\textit{gtr}\leftrightarrow\textit{gte}$ pair. We underline scores that are within a $0.01$ margin, as they might result from rounding noise, stochasticity, and a slightly different experimental setup. [^4] Our results are calculated over multiple runs and a single evaluation set, while vec2vec computes the results of a single run over 8 evaluation sets. Due to the small variance in most experiments, in both methods, we believe that a $0.01$ margin is good enough.
@@
The slightly superior results of mini-vec2vec are very encouraging, but even in pairs where the results are slightly worse or match vec2vec, the main point is that it is achieved for a fraction of the cost. One run of mini-vec2vec is completed on a CPU in less than ten minutes, while vec2vec requires 1–7 days on a GPU, depending on the hardware. Another important point from the point of view of robustness is that even where the vec2vec results are relatively low, mini-vec2vec still performs well, close to perfect performance.
@@
Analyzing the convergence of our algorithm, we observe that despite variation in the approximate matching step (Initial), Refine-1 consistently improves cosine similarity and converges to stationary points with nearly identical cosine similarity values across different random initializations. This reproducible convergence behavior suggests the optimization landscape is well-structured. The convergence of Refine-1 appears to reach a stationary point of the implicit nearest-neighbor matching objective. However, applying Refine-2 successfully improves upon this converged solution. The two methods appear to optimize different implicit proxy objectives, and both contribute to reaching a good solution.
@@
gtr 110M T5 gtr 2021
e5 109M BERT e5 2022
gte 109M BERT gte 2023
stella 109M BERT stella 2023
granite 278M RoBERTa granite 2024
@@
6 Discussion
6.1 Choice of Hyperparameters
@@
  • Number of iterations: We find that the alignment is monotonically increasing as the number of iterations grows, unlike some algorithms where increasing information might cause degradation. This makes it a safe bet to increase the number of iterations. Across all cases examined, though, we found convergence to be quite fast. Almost without exception, 100 iterations were enough, by a large margin.
@@
Other hyperparameters are less obvious. Number of clusters in the approximate matching step is likely good in moderation. While increasing it will give us more landmark points for alignment, it can theoretically lead to instabilities because the assignment of very granular centroids can be noisy. Increasing number of neighbors is often beneficial, but it eventually leads to diminishing returns, and might even deteriorate performance when neighborhoods become very large. At moderate scales, it helps to reduce noise, as averaging more neighbors leads to a smoother estimate of the true aligned position. Note that this is more relevant here than in word embeddings, where essentially every point has (more or less) one ground-truth translation, and it makes sense to rely on a single neighbor. Increasing the number of neighbors to very large numbers might lead to over-smoothing, which may harm performance. Finally, increasing $\alpha$ leads to higher cosine scores, but they don’t necessarily translate to better rank and top-1 scores. $\alpha$ is implicitly responsible for the distance of the matrix from being orthogonal. Additionally, increasing $\alpha$ makes convergence slightly slower.
@@
6.2 Failure Modes
6.3 Alternative Approaches
Through anecdotal experiments, we have found that it might be possible to sidestep the use of relative representations in the approximate matching step. Instead, we can perform Procrustes Analysis directly on the concatenated aligned centroids. It seems to do slightly worse in the approximate matching step, but it converges normally after. We emphasize, as noted above, that we haven’t tested this thoroughly to be certain it generalizes. For future reference, we call this variant mini-vec2vec without relative representations”, but catchier names can be used if meaning is clear.
7 Related Work
The study of universality has its roots in several works. Representation alignment is well-studied in the supervised setup. lenc2014stitching and bansal2021revisiting have demonstrated that linear transformations are often sufficient for aligning neural network representations through their stitching approach. colah2015visualizing studied the equivalence of neural networks in relative terms. moschella2022relative have used these insights to encode a universal space where data points are represented relative to anchor points, utilizing it for zero-shot stitching. Similarly, dar2023analyzing have demonstrated zero-shot stitching using a method equivalent to relative representations with token embeddings as anchors. huh2024platonic have aggregated and unified previous work, under the name Platonic Representation Hypothesis.
The task of unsupervised alignment of embeddings also has a long history in word embeddings (conneau2017word; chen-cardie-2018-unsupervised; grave2018unsupervisedalignmentembeddingswasserstein) and machine translation (lample2018unsupervisedmachinetranslationusing; artetxe2018unsupervisedneuralmachinetranslation). Most of the methods rely on training a GAN, and they often require a small seed of matched pairs in order to work. hoshen-wolf-2018-non avoid the adversarial setup and use a three-step algorithm similar to ours. Their approach, especially the approximate matching step, which relied on PCA, was insufficient for our purposes.
8 Conclusion
We have presented mini-vec2vec, a linear approach to unsupervised embedding alignment that achieves competitive performance while offering substantial advantages in computational efficiency, training stability, and interpretability. Our method demonstrates that adversarial training may be unnecessary for reliable embedding alignment, and identifying structural correspondences is possible in other ways. The success of our approach provides another strong empirical evidence for the Universal Geometry hypothesis and suggests that the geometric regularities in learned representations are more robust and accessible than previously assumed. By decomposing the alignment problem into classical optimization components, we achieve both theoretical clarity and practical efficiency.
The efficiency and accessibility of our method can help democratize research on embedding alignment and cross-modal representation learning. By reducing computational barriers, we enable broader participation in this research area and facilitate deployment in resource-constrained environments. From a security perspective, efficient alignment methods raise important considerations about data privacy and intellectual property protection in embedding spaces. If alignment can be performed quickly and reliably, it may become easier to leak data in the scenario described in jha2025vec2vec.
Acknowledgements

### v4

@@ 3.4 Iterative Refinement
This refinement strategy proceeds by transforming embeddings from space $A$ to $B$ using $\mathbf{W}$, averaging their nearest neighbors in space $B$, and obtaining a new orthogonal transformation with Procrustes analysis – similar to the process described in the Initial Transformation stage (Section 3.3), with the crucial difference that now the similarity is computed in the ambient space $B$ rather than relative space. This technique is also present in Hoshen and Wolf (2018), and it is often called Iterative Closest Point (ICP; Besl and McKay, 1992). A minor difference is that we average multiple points rather than relying on a single neighbor. As mentioned above, this is due to our no-overlap setup. Since ICP is cheap, we run it for many iterations (50–100), using subsampling to boost computational efficiency further.
@@
This strategy aims to improve the large-scale matching between the spaces. We begin by clustering the embeddings in space $A$. We then apply $\mathbf{W}$ to the cluster centroids and cluster the $B$ embeddings, where the clustering algorithm is initialized with the transformed $A$ centroids as the initial centroids. The transformed clusters should be close to a set of clusters in the new space. We therefore expect the clustering algorithm to make only minor adjustments, correcting biases in the transformation and moving the centroids to their “right positions” in space $B$. We expect the pseudo-parallel pairs based on cluster assignments to be more robust than individual nearest neighbors used in Refine-1. Similar to the first refinement strategy, we use exponential smoothing. Crucially, we find that one iteration of Refine-2 improves the alignment obtained by Refine-1. Two or more lead to a slight deterioration of the alignment. We conjecture that the ICP procedure might have introduced systematic bias that requires correction, but after this correction, the clustering-based refinement does not contribute, and only dilutes the signal.
@@ 4 Experiments
We replicate the main experiment from Jha et al. (2025). We run our algorithm for all pairs of encoders used in Jha et al. (2025). Details about the encoders are provided in Table 2 in the Appendix. Experiments use k-means clustering from scikit-learn (Pedregosa et al., 2011) and QAP is solved with the scipy (Virtanen et al., 2020) implementation of 2-OPT. We re-run the 2-OPT alignment algorithm 30 times and choose the permutation with the highest alignment score. This helps improve the robustness of the alignment, as a single 2-OPT is often not enough.
We take a random subset of the Natural Questions dataset (Kwiatkowski et al., 2019) of size $60,000$, compute sentence embeddings with all encoders, and leave $8192$ sentences for evaluation. We split the remaining embeddings equally between the source encoder (space $A$) and target encoder (space $B$), so there is no overlap between the source and target sentences. All experiments are run on a Colab notebook with a CPU runtime and repeated three times for each pair to demonstrate the method’s stability. We compare our results to the numbers reported in Jha et al. (2025).
@@
We have found that the method is extremely robust to hyperparameter choices, as long as they are in a reasonable range. For the paper, we made an economical choice of hyperparameter values to shave time for the large number of runs, but we found the choice to be sufficiently robust even still. We use $s=30$ runs of relative representation alignment with $c=20$ clusters. In the Initial Transformation step, we use $k=50$ neighbors. We set the sampling rate in Refine-1 to $n_{s}=10,000$ and utilize $k^{\prime}=50$ neighbors; In Refine-2, we use $c^{\prime}=500$ clusters. For exponential smoothing, we use $\alpha=0.5$. The full hyperparameter configuration used in the experiments is also documented in the code linked. For the $\textit{gtr}\leftrightarrow\textit{gte}$ pair, we increase the number of clusters to $c=30$. We find that for this pair, the algorithm often converges to a sub-optimal solution in the $c=20$ regime. A detailed discussion of observations about hyperparameter choice can be found in Appendix B.
@@ Results
From the results in Table 1, we observe that our approach matches or exceeds vec2vec across (almost) the entire table. The improvement is most noticeable in the rank metric. Importantly, the method is extremely robust as is evidenced by the small standard deviation, already by the end of the first refinement step. Unlike vec2vec, our method does not collapse on the $\textit{e5}\leftrightarrow\textit{gtr}$ pair or the $\textit{gtr}\leftrightarrow\textit{gte}$ pair. We underline scores that are within a $0.01$ margin, as they might result from rounding noise, stochasticity, and a slightly different experimental setup. [^1] Our results are calculated over multiple runs and a single evaluation set, while vec2vec computes the results of a single run over 8 evaluation sets. Due to the small variance in most experiments, in both methods, we believe that a $0.01$ margin is good enough.
@@
The slightly superior results of mini-vec2vec are very encouraging, but even in pairs where the results are slightly worse or match vec2vec, the main point is that this is achieved for a fraction of the cost. One run of mini-vec2vec is completed on a CPU in less than ten minutes, while vec2vec requires 1–7 days on a GPU, depending on the hardware. Another important point from the point of view of robustness is that even where the vec2vec results are relatively low, mini-vec2vec still performs well, close to perfect performance. Analyzing the convergence of our algorithm, we observe that despite variation in the Initial Transformation step (Initial), Refine-1 consistently improves cosine similarity and converges to stationary points with nearly identical cosine similarity values across different initializations. This reproducible convergence behavior suggests the optimization landscape is well-structured. The convergence of Refine-1 appears to reach a stationary point of the implicit nearest-neighbor matching objective. Refine-2 successfully improves upon this converged solution, which might indicate that the converged solution has some systematic bias.
5 Discussion
Failure Modes
Main Takeaways
We would like to emphasize the subtle points that make the method work. While word alignment methods weren’t suitable in this setup, they are likely not completely irrelevant. The main drivers of success are stable landmarks and tractable matching. We chose to focus on clusters as these landmarks. However, it is not the only option. A preliminary study (Yamagiwa et al., 2023) on word and image embeddings has found that Independent Component Analysis (ICA) can identify shared interpretable directions between embedding spaces. We don’t preclude the possibility that they could have been used instead. The important part is to find anchors that are consistent despite zero overlap. This gives us a match between sets, and leaves us with the task of aligning the individual anchors. The fact that the number of anchors is relatively small makes the process very fast, but also makes the problem simpler. It is easier to find near-optimal solutions when the number of samples is not large, while the problem quickly gets out of hand due to its combinatorial nature. We used QAP for matching, which is closely related to Optimal Transport. The small scale does more than speed up the process, but likely also improves the performance of the algorithm. We believe this might have worked with Gromov-Wasserstein and other OT algorithms as well, as long as the same principles were kept. Last but not least, robustness to noise is another crucial driver of success. While Hoshen and Wolf (2018) re-initialize their pipeline (500 times, which is very expensive, though parallelizable) to fight noise and arbitrariness, we use many noisy assignments in one run, addressing the problem without multiple runs. Other tricks, not necessarily common in literature, such as exponentially-smoothed orthogonal matrices, cluster-based matching, and others, have empirically contributed to boosting performance significantly, but even without them performance was decent.
6 Related Work
Universality
The study of universality has its roots in several works. Representation alignment is well-studied in the supervised setup. Lenc and Vedaldi (2015) and Bansal et al. (2021) have demonstrated that linear transformations are often sufficient for aligning neural network representations in a procedure they called stitching. Mikolov et al. (2013) have identified that word embeddings from different languages converge to similar geometries and used it to align them with a learned linear transformation from a seed dictionary. Olah (2015) visualized representations from different convolutional neural networks using relative representations. Moschella et al. (2022) have used these insights to transition between representation spaces through a universal space, utilizing it for zero-shot stitching. Concurrently, Dar et al. (2023) have demonstrated zero-shot stitching using (a method equivalent to) relative representations with token embedding anchors. Huh et al. (2024) have aggregated and unified previous work under the name Platonic Representation Hypothesis.
Unsupervised Alignment
The task of unsupervised alignment of embeddings, too, has a long history in word embeddings (Conneau et al., 2017; Chen and Cardie, 2018; Zhang et al., 2017) and machine translation (Lample et al., 2017). These early methods relied on adversarial training. Artetxe et al. have started a line of supervised and semi-supervised alignment methods relying on an iterative refinement process (Artetxe et al., 2016, 2017), which culminated in VecMap (Artetxe et al., 2018), an unsupervised method that relied on the distributional features of words. Another unsupervised work (Hoshen and Wolf, 2018) used a three-step algorithm not unlike our own, but it required hundreds of re-initializations and used word frequencies. The works of Alvarez-Melis and Jaakkola (2018) and Grave et al. (2019) have utilized variants of optimal transport to align word embeddings. They require the existence of a match between points, at least approximately, in order to work. Works on unsupervised alignment in other domains, for example, image or cross-modal image-text alignment, have mostly focused on semi-supervised setups and overlapping datasets (Schnaus et al., 2025; Maniparambil et al., 2024), which inspired our initial attempts.
7 Conclusion
We have presented mini-vec2vec, a linear approach to unsupervised embedding alignment that achieves competitive performance while offering substantial advantages in computational efficiency, training stability, and interpretability. Our method demonstrates that adversarial training may be unnecessary for reliable embedding alignment, and identifying structural correspondences is possible in other ways. The success of our approach provides another strong empirical evidence for the Universal Geometry hypothesis and suggests that the geometric regularities in learned representations are more robust and accessible than previously assumed. By decomposing the alignment problem into classical optimization components, we achieve both theoretical clarity and practical efficiency. The efficiency and accessibility of our method can help democratize research on embedding alignment and cross-modal representation learning. By reducing computational barriers, we enable broader participation in this research area and facilitate deployment in resource-constrained environments. From a security perspective, efficient alignment methods raise important considerations about data privacy and intellectual property protection in embedding spaces. If alignment can be performed quickly and reliably, it may become easier to leak data in the scenario described in Jha et al. (2025).
Acknowledgements
Appendix A Additional Assets
@@
gtr 110M T5 Ni et al. 2021
e5 109M BERT Wang et al. 2022
gte 109M BERT Li et al. 2023
stella 109M BERT Zhang et al. 2023
granite 278M RoBERTa Awasthy et al. 2024
@@
Appendix B Choice of Hyperparameters
@@
  • Number of iterations: We find that the alignment is monotonically increasing as the number of iterations grows, unlike some algorithms where increasing information might cause degradation. This makes it a safe bet to increase the number of iterations. Across all cases examined, though, we found convergence to be quite fast. Almost without exception, 100 iterations were enough by a large margin.
@@
Other hyperparameters are less obvious. Number of clusters in the approximate matching step is likely good in moderation. While increasing it will give us more landmark points for alignment, it can theoretically lead to instabilities because the assignment of very granular centroids can be noisy. Increasing number of neighbors is often beneficial, but it eventually leads to diminishing returns, and might even deteriorate performance when neighborhoods become very large. At moderate scales, it helps to reduce noise, as averaging more neighbors leads to a smoother estimate of the true aligned position. Note that this is more relevant here than in word embeddings, where essentially every point has (more or less) one ground-truth translation and it makes sense to rely on a single neighbor. Increasing the number of neighbors to very large numbers might lead to over-smoothing, which may harm performance. Finally, increasing the exponential smoothing hyperparameter $\alpha$ leads to higher cosine scores, but they don’t necessarily translate to better rank and top-1 scores. $\alpha$ is implicitly responsible for the distance of the matrix from being orthogonal. Additionally, increasing $\alpha$ makes convergence slightly slower.